. Spherically symmetric charge distribution. Fig. 3-15 shows a spherical distribution of charge of radius R. The charge density  at any points depends only on the distance of the point from the center and not on the direction, a condition called spherical symmetry. Find an expression for E for points (a) outside and (b) inside the charge distribution. Note that the object in Fig. 3-15 cannot be a conductor. 


1
Expert's answer
2021-05-31T15:38:08-0400

Solution.

"\rho _0" ;

"R;"

"r;"

"a) E_{out}=\dfrac{1}{4\pi\epsilon_0}\dfrac{q_{tot}}{R^2};"

"q=\int\rho_0dV=\rho_0(\dfrac{4}{3}\pi R^3);"

"E_{out}=\dfrac{1}{4\pi \epsilon_0}\dfrac{\rho_0(\dfrac{4}{3}\pi R^3)}{R^2}=\dfrac{\rho_0 R}{3\epsilon_0};"

"b)" "E_{in}=\dfrac{1}{4\pi\epsilon_0}\dfrac{q_{within}}{r^2};"

"q=\int\rho_0dV=\rho_0(\dfrac{4}{3}\pi r^3);"

"E_{in}=\dfrac{1}{4\pi \epsilon_0}\dfrac{\rho_0(\dfrac{4}{3}\pi r^3)}{r^2}=\dfrac{\rho_0 r}{3\epsilon_0};"

Answer: "a)E_{out}=\dfrac{\rho_0 R}{3\epsilon_0};"

"b)E_{in}=\dfrac{\rho_0 r}{3\epsilon_0}."